# Heavy-element damage seeding in proteins under XFEL illumination

### v4

@@ Abstract
Serial femtosecond X-ray crystallography (SFX) captures the structure and dynamics of biological macromolecules at high spatial and temporal resolutions. The ultrashort pulse produced by an X-ray free electron laser (XFEL) ‘outruns’ much of the radiation damage that impairs conventional crystallography. However, the rapid onset of ‘electronic damage’ due to electron ionization limits this benefit. Here, we distinguish the influence of different atomic species on the electronic damage to protein crystals in SFX by employing a plasma code that tracks the unbound electrons as a continuous energy distribution. The simulations show that trace quantities of heavy atoms (10) contribute a substantial proportion of global radiation damage by inducing electron ionization processes that are highly damaging on the XFEL pulse timescale. Native sulfur atoms have a considerable effect on damage-induced noise. Likewise, proteins suffer less damage when the crystal solvent contains lower concentrations of heavy elements. The proximity of the X-ray energy to a heavy-element absorption edge has a strong and non-trivial influence on the damage induced in these models. In one representative protein (lysozyme substituted with selenomethionine), adjusting the X-ray energy from 14 keV to 12 keV is as effective at reducing the ionization of light atoms as compressing the pulse width from 50 fs to 5 fs. These results indicate that relatively small quantities of heavy elements can substantially affect global radiation damage in the XFEL regime.
@@ I Introduction
Serial femtosecond crystallography (SFX) is a potentially revolutionary structural determination tool for biology that overcomes key challenges faced by conventional X-ray crystallography [1]. Using X-ray free electron lasers (XFELs) to serially illuminate small crystals with ultrabright, femtosecond pulses of radiation, a molecule’s structural signal is captured so swiftly that the atomic nuclei are effectively frozen in place [2, 3, 4]. This approach not only mitigates structural damage, allowing for the use of much higher fluences, but also facilitates time-resolved crystallography. As such, SFX can be applied to capture molecular movies of rapid biological processes, including photoactivated dynamics [5, 6, 7, 8, 9, 10, 11] and enzyme catalysis [12, 13, 14].
@@
In practice, however, ultrafast electronic processes disrupt the simple diffract-and-destroy picture – scattering signals seen in experiments generally remain dose-dependent even at the practical minimum pulse widths achievable ( ) [26, 27, 17, 28]. In particular, the electrons ejected through photoionization and subsequent Auger decays initiate electron impact ionization (EII) cascades that will go on to become the main cause of ionization well before the end of the pulse [3, 29]. This ‘electronic damage’ has complex, non-linear effects on the scattered wavefield [26, 30, 31, 32], and various theoretical treatments have been proposed to model the underlying ultrafast damage processes [32, 30, 33, 31, 34, 29]. Efficient use of limited XFEL beamtime [21, 35] therefore often depends on identifying experimental parameters that minimize electronic damage [36].
@@
Heavy atoms ( in this context) are ubiquitous in protein crystals. Though they typically comprise less than 1% of the illuminated atoms, they play a key role in the biochemistry of metalloproteins. Additionally, their unusual scattering behavior has various applications to X-ray crystallography. Anomalous phasing, a notable example, has been successfully extended to SFX with microcrystals using sulfur native to proteins [37, 38, 39], or with much heavier atoms deliberately added for a stronger signal [40, 41, 42] (as in Fig. 1). The rapid ionization of heavy atoms under high fluences induces a strong time dependence in their atomic scattering factors. This behavior is central to proposed methods of high-intensity radiation-damage-induced phasing (HI-RIP) [43, 44, 45, 46, 47, 48, 27, 49, 3].
@@
A large part of the existing body of work dedicated to simulating atomic dynamics under XFEL illumination follows either the Monte Carlo molecular dynamics (MD) paradigm [28, 50, 51, 52] or a plasma approach [53, 54, 33]. These frameworks are in some sense complementary; the former provides a detailed picture of damage on the local scale, while the latter focuses on global statistics [49]. In the MD case, individual ions and electrons are treated as classical particles and tracked through space. Such modeling is too computationally demanding for system sizes over 100–1000 atoms, so cannot simulate typically sized proteins in full [27, 55, 32, 34]. In contrast, models that incorporate a zero-dimensional plasma code may capably simulate much larger (or arbitrary) systems [53, 33, 56, 49], mitigating possible difficulties in capturing the effect of trace elements.
@@
Several plasma codes that treat the free electrons out of thermal equilibrium have been applied in studies of single-element targets [57, 34, 54, 58, 59]. However, all plasma codes used to simulate the damage sustained by biomolecules model the free electrons in local thermal equilibrium (LTE) with a Maxwellian energy distribution. This treatment may miss important effects – existing work on single-element, solid-density targets quotes thermalisation times in picoseconds [60, 61, 57, 62, 34, 58].
@@
In this work, we present a plasma physics code with a non-LTE continuum model of the electron energy distribution and, using this code, examine electronic damage in protein crystals containing heavy elements without assuming instantaneous thermalization. We combine a custom-built frozen-shell Hartree-Fock code [63] with a non-standard B-spline approach to solving the Boltzmann equation for the time-dependent non-equilibrium energy distribution of the free electrons. The details of this framework are given in Sec. II. We apply the model to compare how light and heavy atoms influence the ionization dynamics of biological matter in Sec. III. Sec. IV identifies how trace quantities of heavy elements in biomolecular targets give rise to species-dependent high-damage regions of the pulse parameter space. After reporting on how the presence and composition of the solvent may affect these dynamics in Sec. V, we go on to identify new possible avenues for mitigating the impact of damage on structure determination in Sec. VI.
@@ II Simulation Method
In a typical XFEL experiment, it is change of ionic state, rather than nuclear motion, that makes up the bulk of the radiation damage [32, 2, 27]. It follows that a simulation of an XFEL target that correctly estimates the time-dependent distribution of ion states will capture the most important part of the damage dynamics. We therefore approach the XFEL dynamics problem from the plasma physics perspective previously used in the study of metal plasmas [60, 61, 57, 58, 59].
This work uses a new version of the collisional-radiative plasma physics code AC4DC [33], which has been extended to couple the non-local thermal equilibrium (non-LTE) electron gas to the ion plasma. The code is optimized to model the plasma dynamics of low- matter in the XFEL regime, treating the constituent atoms as independent ions coupled to a bath of free electrons and the driving XFEL photons.
@@
The model couples the time-dependent energy distribution of the free electrons, , to the population of possible ionic states, , through the processes of photoionization, Auger and fluorescent decay, electron impact ionization (EII), three-body recombination (TBR), and pairwise Coulomb electron-electron interactions. These processes are summarized in Fig. 2. Atomic parameters are calculated in the radially-averaged Hartree-Fock approximation [33, 63]. EII and TBR are approximated using the well-established binary-encounter dipole model of Kim and Rudd [67]. The equations of motion, obtained from radially averaging the Boltzmann equation, then read
@@
where and represent the couplings to processes affecting and , respectively (illustrated in Fig. 2). Details for the calculations of the atomic cross-sections, , are given in Ref. [63]. Explicit expressions for are given in Appendix A.
@@
We discretize the free electron distribution using an adaptive-grid spline expansion, in which is expanded in piecewise-polynomial B-splines [68]. These basis functions have compact support, allowing for efficient computations of the second- and third-order tensors without sacrificing differentiability of . As the simulation progresses, the spline grid is dynamically transformed to have an increased knot density in regions with complex energy-space structure, such as the vicinities of the thermal and photoelectron peaks (see Appendix A for further details). This adaptive grid allowed the code to perform full dynamical non-equilibrium plasma simulations of lysozyme in 1 hour on a contemporary desktop.
@@
The code achieved excellent agreement with published Monte-Carlo simulations by ddcMD of amorphous carbon [34] for the evolution of the ionic states and free electrons, and predicted similar ionic state dynamics to simulations by XMDYN of a glycine crystal [50, 32]. Somewhat surprisingly, the code also sees good agreement with the particle-in-cell DFT code PICLS [57] when modeling aluminium plasma for both the ion and free electron populations.
@@
We first study the impact of heavy atoms in a representative model system – gadolinium-derivative hen egg-white lysozyme (lysozyme.Gd, derived from PDB entry 4ET8 [69, 17]) – subjected to a 15 fs FWHM Gaussian pulse of fluence ph -2. This protein derivative contains 2 Gd3+ ions and 10 S atoms in each asymmetric unit of 1000 atoms. Three ‘toy’ variants of this system were simulated: lysozyme.Gd, non-derivative (Gd-free) lysozyme, and lysozyme with sulfur atoms substituted for nitrogen (the ‘light-atom control’). By contrasting the damage in these three materials, the effect of the heavy atoms on the dynamics may be inferred. The presence of water was neglected in these simulations; each system was modeled as a continuous, homogeneous material with the atomic composition of the protein – see Sec. V for an examination of the influence of the solvent within the protein crystal.
Figure 3(a) shows snapshots of the electron distributions of the systems during the pulse. Though S and Gd hold only a small fraction of the electrons in the lysozyme.Gd system, it may be readily seen that their photoelectrons hold a substantial fraction of the free electrons’ energy, comparable to that held by the light-atom photoelectrons. This is a direct consequence of the dependence of photoabsorption cross-sections on atomic number [70]. The photoelectron peaks remain sharp for a substantial part of the pulse (up to ), emphasizing the necessity of the non-LTE treatment.
The inset plots in Fig. 3(a) show lysozyme.Gd with approximately twice the population of thermal electrons as the light-atom control. The vast majority of the additional electrons come from extra secondary ionization events in the light atoms. This is reflected by Fig. 3(b), which shows that the carbon 2s and 2p orbitals deplete much more rapidly when S and Gd are present in the system.
Figure 4 shows the heavy elements in lysozyme.Gd substantially elevate the degradation of the diffraction pattern due to the additional electronic damage suffered by the light atoms (see Fig. S3 for the effect on the atomic form factors). The diffraction pattern was computed by integrating snapshots of lysozyme’s light atom structure, with each atom’s electronic configuration, and thus scattering profile, stochastically chosen with weightings according to the bound state probability distribution of the plasma simulation (see Appendix A). We considered a single-protein diffraction pattern to highlight the impact of damage on the underlying signal without the need for assumptions on crystal size, morphology, and quality.
As is standard for studies of radiation damage [2, 29, 30, 32], this work employs an factor as a metric for the severity of the radiation damage – specifically, the weighted mean error between the ideal and damaged scattering patterns . While directly comparing this measure with experimental factors would be misleading, it may be reasonably used to predict the relative contribution of damage to experimentally analogous noise in initially identical structures (a constrained usage compared to prior works, see Appendix B). The presence of gadolinium and sulfur increases by 59% (relative to the light atom control), indicating their contributions to damage-induced noise are substantial.
Simulations performed with RADDOSE-XFEL, a Monte Carlo code that does not include high-density effects like TBR or the depletion of the ionic states, also predicted a strong impact by gadolinium and sulfur on the light atoms’ ionization. These simulations demonstrated that excluding later ionization is key to the effect: calculating the absorbed doses with RADDOSE-3D, which assumes EII to be instantaneous, grossly underestimated the relative contribution of heavy atoms.
@@ III.1 Comparing the ionization behavior of light and heavy elements
The primary electron production rate rises substantially with higher . This is partly due to the scaling of the photoabsorption cross-section. It also results from the fall in decay transition lifetimes – from 10 fs in carbon to near or below 1 fs for elements heavier than sulfur [19, 71] – which both mitigates frustrated absorption resulting from core holes [72] and shortens the delay between photoionization and Auger electron emission. This suggests that on short timescales, the influence of heavy atoms on electronic damage is mainly distinguished by their primary ionization behavior. It can be seen how this could make heavy elements significant sources of damage by picturing a heavy atom within a mass of light atoms exposed to an XFEL pulse. The heavy atom undergoes rapid primary ionization at early times, seeding many electron cascades that each impact and ionize many light atoms over 3–20 fs [3].
To verify this picture, AC4DC was used to simulate the electron dynamics of a homogeneous ‘Fe-doped’ target containing 1 Fe2+ ion for every 99 light atoms under four models:
a)
Primary ionization only.
b)
All ionization in CNO only.
c)
All ionization in CNO and primary ionization in Fe.
d)
All ionization processes.
The time-dependent ionic states of these systems under a 15 fs pulse with a square temporal profile are visualized in Fig. 5. The pulse was simulated at a fluence on the upper end of modern capabilities ( 10 keV ph-2), where the influence of secondary ionization is comparatively weak. Contrasting models a) and b) with the complete dynamics of model d), it is clear that both light-atom secondary ionization and the presence of the Fe ions are essential parts of the dynamics. The similarity of models c) and d) shows that the influence of the Fe ions almost entirely originates with their primary ionization processes.
The rapid ejection of electrons from the K-shells of Fe ions is sustained by the electrons in higher orbitals, which fill core holes on a subfemtosecond timescale through Auger decay and, to a lesser extent, fluorescence. As a result, the 1s orbital remains almost full until the higher shells are stripped. This Auger cycling (a mechanism previously identified by Refs. [52] and [73]) implies the early-time photoionization rate of the Fe ions scales almost linearly with photon intensity.
From this perspective, it can be understood why the heavy elements in lysozyme and lysozyme.Gd have a disproportionate influence on electronic damage. Although the electrons ejected from light atoms make up the overwhelming majority of the free electron population, most of these electrons are ejected by EII cascades, of which a substantial fraction are seeded by the primary ionization of S and Gd atoms (see Fig. 6). The amount of damage caused by a cascade will depend on the energy of the instigating primary electron, so the order 1 keV separations between the CNO, S, and Gd photoelectron and Auger emission energies observable in Fig. 3(a) lead to further differences in how they affect the target’s secondary ionization. See Sec. IV.1 for further details. It should be noted that in the lysozyme.Gd simulation, the ionizable shells of Gd (M, N, and O) were almost completely stripped of electrons by ; in reality, the highly charged Gd ions would be replenished by bound electrons transported from neighbouring atoms [73], likely leading to a stronger damage seeding effect.
Finally, we remark on the importance of the pulse profile on these dynamics. Previous works have generally modeled the temporal intensity profile as either square [34, 57] or Gaussian [57, 58]. The ionization dynamics proved to be sensitive to this choice (see Fig. S1). Repeating the lysozyme.Gd simulation (Fig. 3) with a square pulse of the same fluence, energy, and FWHM showed to be 30% higher in the Gaussian case. This can be attributed to the outsized effect of earlier EII cascades [74], induced by the leading tail of the pulse. Such cascades have a long period to ionize the target before the bulk of the elastic X-ray scattering.

### v5

@@ Abstract
Serial femtosecond X-ray crystallography (SFX) captures the structure and dynamics of biological macromolecules at high spatial and temporal resolutions. The ultrashort pulse produced by an X-ray free electron laser (XFEL) ‘outruns’ much of the radiation damage that impairs conventional crystallography. However, the rapid onset of ‘electronic damage’ due to ionization limits this benefit. Here, we distinguish the influence of different atomic species on the ionization of protein crystals by employing a plasma code that tracks the unbound electrons as a continuous energy distribution. The simulations show that trace quantities of heavy atoms () contribute a substantial proportion of global radiation damage by rapidly seeding electron ionization cascades. In a typical protein crystal, sulfur atoms and solvated salts induce a substantial fraction of light-atom ionization. In further modeling of various targets, global ionization peaks at photon energies roughly 2 keV above inner-shell absorption edges, as sub-2 keV electrons initiate ionization cascades that are briefer than the XFEL pulse. These results indicate that relatively small quantities of heavy elements can substantially affect global radiation damage in XFEL experiments.
@@ I Introduction
Serial femtosecond crystallography (SFX) is a potentially revolutionary structural determination tool for biology that overcomes key challenges faced by conventional X-ray crystallography [1]. Using X-ray free electron lasers (XFELs) to serially illuminate small crystals with ultrabright, femtosecond pulses of radiation, the structural signal of a molecule is captured so swiftly that the atomic nuclei are effectively frozen in place [2, 3, 4]. This approach not only mitigates structural damage, allowing for the use of much higher fluences, but also facilitates time-resolved crystallography. As such, SFX can be applied to capture molecular movies of rapid biological processes, including photoactivated dynamics [5, 6, 7, 8, 9, 10, 11] and enzyme catalysis [12, 13, 14].
@@
In practice, ultrafast electronic processes disrupt the simple diffract-and-destroy picture – scattering signals seen in experiments generally remain dose-dependent even at the practical minimum pulse widths achievable ( ) [26, 27, 17, 28]. In particular, the electrons ejected through photoionization and subsequent Auger decays initiate electron impact ionization (EII) cascades that will go on to become the main cause of ionization well before the end of the pulse [29, 3, 30]. This ‘electronic damage’ has complex, non-linear effects on the scattered wavefield [26, 31, 32, 33], and various theoretical treatments have been proposed to model the underlying ultrafast damage processes [33, 31, 34, 32, 35, 30]. Efficient use of limited XFEL beamtime [21, 36] therefore often depends on identifying experimental parameters that minimize electronic damage [37].
@@
Heavy atoms ( in this context) are ubiquitous in protein crystals. Though they typically comprise less than 1% of the illuminated atoms, they play a key role in the biochemistry of metalloproteins. Additionally, their unusual scattering behavior has various applications to X-ray crystallography. Anomalous phasing, a notable example, has been successfully extended to SFX with microcrystals using sulfur native to proteins [38, 39, 40], or with much heavier atoms deliberately added for a stronger signal [41, 42, 43]. The rapid ionization of heavy atoms under high fluences induces a strong time dependence in their atomic scattering factors. This behavior is central to proposed methods of high-intensity radiation-damage-induced phasing (HI-RIP) [44, 45, 46, 47, 48, 49, 27, 50, 3].
@@
A large part of the existing body of work dedicated to simulating atomic dynamics under XFEL illumination follows either the Monte Carlo molecular dynamics (MD) paradigm [28, 51, 52, 53] or a plasma approach [54, 55, 34, 56]. These frameworks are in some senses complementary; the former provides a detailed picture of damage on the local scale, while the latter focuses on global statistics [50]. In the MD case, individual ions and electrons are treated as classical particles and tracked through space. Such modeling is too computationally demanding for system sizes over 100–1000 atoms, so cannot simulate typically sized proteins in full [27, 57, 33, 35]. In contrast, models that incorporate a zero-dimensional plasma code may capably simulate much larger (or arbitrary) systems [54, 34, 58, 50], mitigating possible difficulties in capturing the effect of trace elements.
@@
Several plasma codes that treat the free electrons as non-thermal (i.e. non-Maxwellian) have been applied in studies of single-element, solid-density targets [59, 35, 55, 60, 56, 61, 62]. However, plasma codes used in prior studies to simulate the damage sustained by biomolecules all approximate the free-electron energy distribution with a Maxwellian, under an assumption of instantaneous thermalization. This treatment may miss important effects – existing work on single-element, solid-density targets quotes thermalisation times in picoseconds [63, 64, 59, 65, 35, 61].
@@
In this work, we employ a non-thermal plasma physics code to examine electronic damage in protein crystals containing heavy elements without assuming instantaneous thermalization. We combine a custom-built frozen-shell Hartree-Fock code [66] with a non-standard B-spline approach to solving the Boltzmann equation for the time-dependent non-equilibrium energy distribution of the free electrons. The details of this framework are given in Sec. II. We apply the model to compare how light (C, N, O) and heavy atoms influence the ionization dynamics of biological matter in Sec. III. We go on to identify how trace quantities of heavy elements in biomolecular targets give rise to species-dependent high-damage regions of the pulse parameter space in Sec. IV. We discuss new avenues for mitigating radiation damage implied by these findings in Sec. V.
@@ II Simulation Method
This study introduces a new non-thermal extension of the atomistic collisional-radiative plasma solver AC4DC described in [66]. While the previous version assumed free electrons to instantaneously thermalize after collisional-ionization, the new treatment allows for a non-Maxwellian free-electron energy distribution of arbitrary functional form. The scope of this study is restricted to the ionization dynamics, neglecting nuclear motion. This focus is motivated by two points: (I) Ionization drives the ionic motion (Coulomb explosion) during and immediately following the XFEL pulse. (II) Studies of XFEL radiation damage consistently show changes in ionic state (electronic damage) make up the bulk of the radiation damage for pulses on the order of 10 fs [33, 2, 27].
@@
The model couples the free-electron energy distribution, , to the population of possible ionic states, , through the processes of photoionization, Auger and fluorescent decay, electron impact ionization (EII), three-body recombination (TBR), and pairwise Coulomb electron-electron interactions (EE). These processes are summarized in Fig. 1. Atomic parameters are calculated in the radially-averaged Hartree-Fock approximation [34, 66]. EII and TBR are approximated using the well-established binary-encounter dipole model of Kim and Rudd [67]. The equations of motion, obtained from radially averaging the Boltzmann equation, then read
@@
where and denote electron configurations, and and represent the couplings to processes affecting and , respectively (illustrated in Fig. 1). Details for the calculations of the atomic cross-sections, , are given in Ref. [66]. In terms of separate processes, has the form
The notation adopted in Eq. (3) is such that each is multilinear in all arguments. Explicit expressions for these source terms and collision kernels are given in Appendix A.
@@
The new code tracks an arbitrary non-Maxwellian distribution by expanding with piecewise-polynomial B-splines [68]. These basis functions are non-zero only for a small region of energy, allowing for an efficient sparse internal representation of without sacrificing the differentiability of . As the simulation progresses, the spline grid is dynamically adapted to have increased density across the energy range spanning the bulk of the thermal electrons and in the vicinity of the main primary ionization peaks (see App. A for further details). This adaptive grid allowed the code to perform full dynamical non-thermal plasma simulations of lysozyme in under an hour on a contemporary desktop.
The total number of modeled atomic configurations for an element scales like the factorial of the total number of subshells of the ground state, making heavier elements very costly to simulate. For computational expedience, shells of elements were modeled with a single angular momentum (i.e. a single energy level). Test simulations showed ionization rates of light and heavy elements were barely affected by this approximation in the regime considered in this work (see Fig. S3).
@@
We quantitatively compared this new version of AC4DC against a number of published results in the literature, which are presented in App. C. The code achieved excellent agreement with non-Maxwellian Monte-Carlo simulations by ddcMD of amorphous carbon [35] for the evolution of the ionic states and free electrons, and predicted similar ionic state dynamics to simulations by XMDYN of a glycine crystal [51, 33]. Somewhat surprisingly, the code also sees good agreement with the particle-in-cell DFT code PICLS [59] when modeling aluminum plasma for both the ion and free-electron populations. AC4DC did not reproduce results for simulations of silicon [55] by a plasma code that uses a similar physical framework to AC4DC, but which fits to a standard grid rather than using adaptive splines (note we found the adaptive grid crucial to achieving convergence, see App. A).
@@
We first study the impact of heavy atoms in a representative model system – gadolinium-derivative hen egg-white lysozyme (lysozyme.Gd, derived from PDB entry 4ET8 [69, 17]) – subjected to a 15 fs FWHM Gaussian pulse of fluence ph -2. The target contains 35.1% (v/v) solvent (10% m/v NaCl, 0.1 M Na acetate, 0.1 M Gd [27]), which we hereafter refer to as the 0.1 M Gd 10% NaCl solvent for brevity. The resulting unit cell, including disordered solvent atoms, has a chemical composition of H13 259C5153N1596O4009S80Gd21Na93Cl87.
Fig. 2(a) shows the free-electron energy distribution of the lysozyme.Gd system at times -10 fs, 0 fs, and +10 fs from the pulse peak. To infer the effect of the heavy atoms on the dynamics, simulations of three ‘toy’ variants of the system are also shown: lysozyme.Gd in 0.1 M Gd solution (H13 846C5143N1596O4300S80Gd21), lysozyme in water (H13 942C5056N1576O4386S80), and a ‘light-atom control’ (H13 942C5056N1656O4386). These have the same solvent density (1.1 g/mol) and protein concentration as the complete system. It can be seen in Fig. 2(a) that heavy elements make significant contributions to the high-energy region of the distribution relative to their small presence. In addition, the inset plots show a substantially larger thermal (low-energy) electron population for the complete target than the light-atom control. The vast majority of the additional thermal electrons come from additional secondary ionization events in the light atoms. This is reflected by Fig. 2(b), which shows that the carbon 2s and 2p orbitals deplete much more rapidly in the complete target.
@@ III.1 Comparing the ionization behavior of light and heavy elements
Fig. 3 shows that the influence of the heavy ions on global ionization almost entirely originates with their primary ionization processes. Modeling that neglects the secondary ionization of heavy atoms (Fig. 3(b)) still produces a nearly identical evolution in the carbon ionic states to those of the complete simulation shown in Fig. 3(c). In other words, the effect of heavy elements on the dynamics is almost entirely due to their primary electron emissions.
The primary electron (combined Auger and photoelectron) production rate rises substantially with higher . This is partly due to the huge scaling of the photoabsorption cross-section ( for the K-shell [70]), and partly due to shorter Auger lifetimes of heavier elements ( fs for elements heavier than sulfur) [19, 71, 72]. Fig. 3(c) shows the M (3n) shells of the Gd ions are nearly full for the entire simulation. The rapid ejection of photoelectrons from the M shell is sustained by the electrons in higher orbitals, which fill core holes on a subfemtosecond timescale through Auger decay and fluorescence. Previous experimental studies have noted such ‘Auger cycling’ significantly elevates primary ionization under XFEL pulses [53, 73], though here it also elevates the number of secondary ionization cascades seeded by Gd.
To quantify the contribution of each element to global ionization in terms of the damage processes that they seed through primary ionization, we construct a measure of the response of the full free-electron energy distribution to a particular source of primary electrons. We introduce auxiliary distributions that partition , i.e. , such that they represent the free-electron density associated with specific primary ionization processes, e.g. C Auger, or O photoionization. Explicitly, we decompose the right-hand side of Eq. (3) as
This choice of decomposition assigns all EII-sourced secondary electrons to the cascade of the impactor (), while the density arising from TBR and EE processes between different cascades is evenly divided between the two.
Fig. 4 shows the aggregate density of the cascades seeded by each atomic-species population. Comparing the light-atom charges (Fig. 4(a)) with the density of the light-atom-seeded cascades (Fig. 4(b)) highlights that the global ionization induced by an atomic species is decoupled from its ionization rate. C, N, and O are ionized at similar rates, largely by EII; however, the O atoms seed a significantly larger fraction of the EII cascades, owing to the scaling of the K-shell photoabsorption cross-section. It can be further seen that the free-electron density seeded by each heavy-element population (Fig. 4(c)) is comparable to that seeded by the far more abundant light-atom populations. Strikingly, the combined density of the Gd-seeded cascades is near or above that of the O-seeded cascades at all times, despite the system containing O and Gd in a ratio of 191:1. Combined, the cascades seeded by heavy elements account for the majority of the freed electrons in the system for the entirety of the pulse.
From this perspective, it can be understood why the heavy elements have a disproportionate influence on the light-atom ionization. The rapid primary ionization of heavy atoms means they seed a significant number of EII cascades in the target despite their trace presence. Since the vast majority of the ionization in the target is due to EII [29, 3], a substantial fraction of ionization in the light-atom structure occurs through cascades seeded by the heavy elements.
Note that the density of the cascades seeded by Gd Auger electrons (Fig. 4(b)) sees negligible growth after as a result of the depletion of the O- and N-shells, observable in Fig. 3(c). By fs, the average Gd ion charge is above +20. In reality, the highly charged Gd ions would be replenished by bound electrons transported from neighboring atoms [73], likely leading to a stronger damage seeding effect.
III.2 Examining unusual Gd charge measurements in experiments
The ionization of Gd ions in lysozyme.Gd under XFEL pulses has previously been investigated by Refs. [27] and [48]. The charge density of the Gd ions is inferred from the integrated electron density around the Gd site relative to a reference region of light atoms, which we refer to as the electron density ratio (EDR). Nominally, this is proportional to
where is the normalized temporal pulse profile, and is the average number of electrons bound to element in the Gd site or light-atom reference region. Assuming the intensity-averaged light-atom ionization is small, the charge gained by Gd can be approximated. Notably, this measure indicated the ionization of Gd to be unexpectedly low in both studies. Fig. 5 compares EDR values obtained from pump-probe experiments by Ref. [27] (using a light-atom reference region of C20N10O10) with predictions from simulations of these experiments by AC4DC using the solvated lysozyme.Gd target. The simulated EDRs accurately reproduce the features of the experimental data, but rescaled by a factor of .
For the four EDR values where the total pulse energy has a range of 0.95–0.99 mJ (35 fs, 37 fs, 62 fs, and 112 fs), the EDR rises with increasing probe delay. We argue this is consistent with the picture of the dynamics described in Sec. III.1, which highlights that secondary ionization – the only ionization during the ‘dead time’ between pump and probe – affects the light atoms far more strongly than the heavy atoms. This implies that the EDR does not gauge the electron density of the heavy atoms but instead serves as a rough indicator of the ratio of secondary ionization to primary ionization. Under this perspective, the EDRs at 0 fs and 102 fs – which correspond to lower total pulse energies – are raised due to the greater significance of secondary ionization at lower fluences.
It is difficult to diagnose the cause of the factor of 2 systematic offset between the predicted and observed EDRs in Fig. 5. A likely contributing factor is the assumption of a homogeneous spatial profile in the AC4DC model. In the experiment, the beam was nominally focused to a FWHM Gaussian, smaller than the lysozyme.Gd nanocrystals [27]. Photons scattering outside the focus will sample a less-ionized region where, repeating the reasoning from earlier, secondary ionization will be relatively strong and the EDR will be higher. Prior work has highlighted that such photons can make up a substantial fraction of detected photons [74, 48, 75], suggesting this correction could be quite strong. An additional cause may be bound-to-bound electron transport, which AC4DC neglects. As alluded to in Sec. III.1, electron transport was previously shown to slow the charge gain of iodine in iodobenzane [73], a molecule with fewer light atoms than the Gd complex. Lastly, it is possible that the experimental measurements of the light-atom reference region did not capture all bound electrons. We note the observed EDRs suggest implausibly high ionization rates: The higher EDR observed by the pulse corresponds to an intensity-averaged light-atom charge of +3.82, ignoring Gd ionization. For reference, the considered work predicted the average light-atom charge to remain below +1.6 at all times.
Using a similar methodology to Ref. [27], Ref. [48] also attempted to measure the charge of Gd in two cases – ‘high fluence’ and ‘low fluence’, with fluences peaking at 7.8 and ph-2 respectively. The work reports only the difference in intensity-averaged Gd charge between these cases, termed the Gd ‘ionization contrast’, as measured by computing the difference in the EDRs and assuming negligible light-atom ionization. A Gd ionization contrast of 8.8–12 electrons per Gd was observed; however, theoretical modeling using the XATOM toolkit predicted that the Gd charge difference ought to be 25 under a spatially homogeneous intensity. We repeated this modeling using AC4DC, finding an even more extreme 33.1. This apparent discrepancy between theory and experiment is substantially reduced by accounting for light-atom ionization lowering the denominator in Eq. (5) (see Fig. S4) – we obtain a Gd ionization contrast of 21.4 after making this correction. Modeling the cascades seeded by the salt (namely Cl) and Gd ions in the crystal was crucial to seeing this level of agreement, due to their significant contribution to light-atom ionization. Without doing so, the simulated Gd ionization contrast only falls to 25.2 (note the actual Gd charge difference remains at 33.1 in either case).